OpenStudy (anonymous):

Mrs. Jennings spent 14 1/6 hours at the gym last month. Mr. Jennings spent 12 5/6 hours at the gym last month. How much more time did Mrs. Jennings spend at the gym than Mr. Jennings? 2/3 1/6 1 1/4 1 1/3

hero (hero):

14 1/6 - 12 5/6 = (14 - 12)(1/6 - 5/6) = 2 (-4/6) = 2 (-2/3) = 6/3 - 2/3 = 4/3 = 1 1/3

hero (hero):

1 1/3 is the answer

OpenStudy (anonymous):

ok thnx!! :)